OpenStudy (anonymous):

A painting is purchased for $500. If the value of the painting doubles every 5 years, then its value is given by the function V(t) = 500 • 2t/5, where t is the number of years since it was purchased and V(t) is its value (in dollars) at that time. What is the value of the painting ten years after its purchase? Answer: A)$2,000 B)$1,400 C)$1,800 D)$1,000 help please:)

OpenStudy (anonymous):

\[V(t)=500\times2^{\frac{t}{5}}\]

OpenStudy (anonymous):

\[V(10)=500\times 2^{\frac{10}{5}}\] \[V(10)=500\times 2^2=500\times 4=2000\]

OpenStudy (anonymous):

Thank you!